OpenStudy (anonymous):

Question 4 (Essay Worth 7 points) [5.05] MA.6.A.3.3 Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b plus one equals thirty Step 1: fifteen times b divided by fifteen plus one equals thirty divided by fifteen Step 2: b plus one equals two Step 3: b plus one minus one equals two minus one Step 4: b equals one

OpenStudy (anonymous):

Translate the sentence into a mathematical equation. Fifteen times b plus one equals thirty. 15 x b + 1 = 30 15b + 1 = 30

OpenStudy (anonymous):

The steps to solve this equation for b involve performing "opposite" operations. The first step would be to subtract 1 from both sides of the equation. 15b + 1 - 1 = 30 - 1 15b = 29 The second step would be to divide both sides by 15. 15b = 29 --- --- 15 15 b = 29/15

OpenStudy (anonymous):

Look through the steps that are in the problem. Do you see any that seem incorrect?

OpenStudy (anonymous):

Step 1 says that you are dividing "15b" by 15 and "30" by 15. This leaves out the +1 in the original equation. Therefore, Step 1 is incorrect because it does not divide each part of the equation by 15.